Use natural logarithms to solve 6e^4x-2=3
1
Expert's answer
2011-02-28T11:18:33-0500
6 e4x - 2 = 3
6e4x = 5
e4x = 5/6
4x = ln (5/6)
x = 1/4 ln (5/6) ≈ - 0.0456